### v2

# Engineering Quantum Reservoirs through Krylov Complexity, Expressivity and Observability

Abstract
This study employs Krylov-based information measures to understand task performance in quantum reservoir computing, a sub-field of quantum machine learning. In our study we show that fidelity and spread complexity can only explain the task performance for short time evolutions of the quantum systems. We then discuss two measures, Krylov expressivity and Krylov observability, and compare them to task performance and the information processing capacity. Our results show that Krylov observability exhibits almost identical behavior to information processing capacity, while being three orders of times faster to compute. In the case when the system is undersampled Krylov observability best captures the behavior of the task performance.
preprint: APS/123-QED
I Introduction
Quantum machine learning (QML) models attempt to exploit the exponential scaling of Hilbert space for computation. A subfield of QML that has garnered significant attention in physics is quantum reservoir computing (QRC). QRC is inspired by classical reservoir computing, where physical systems are used to solve machine learning tasks [1, 2, 3, 4, 5, 6, 7]. QRC research, particularly with the Ising model, has been explored in [8, 9, 10, 11, 12, 13, 14, 15], with initial implementations on IBM’s quantum processor [16, 17, 18, 19]. It has been shown that quantum reservoir computing can utilize noise and dissipation as resources [20, 21, 22]. To address the time complexity problem for time-series tasks, weak measurements [23, 24] and reinitialization schemes [25] have been proposed, reducing complexity and improving performance. Other approaches include reintroducing measured outputs [26, 27] and continuous measurements [28], while [29] utilizes techniques from parameterized quantum circuits to define expressive limits in QRC. However, understanding in quantum machine learning and quantum reservoir computing remains limited. In classical machine learning, this understanding arises from expressivity and explainability measures, which is still an active field and is mostly unresolved in quantum machine learning. A key challenge is comparing quantum systems described by unitaries or Hamiltonians based on their expressivity within the Hilbert space. An ideal information measure would provide insights into task performance while offering a physical interpretation of the system, making it physically interpretable.
In this work we discuss various measures regarding task performance in quantum reservoir computing. We will train four distinct quantum reservoirs on a chaotic time series. As a benchmark the Lorenz63 system is used [30]. We compare fidelity, information processing capacity [31], and spread complexity [32, 33] as established methods. Additionally, we discuss Krylov expressivity and Krylov observability, two measures designed to capture the phase space dimension of the Hilbert space[34, 35]. We demonstrate that Krylov observability and information processing capacity perform best in explaining the behavior of task performance.
Spread and operator complexity have successfully captured various effects of time evolution in relation to spin chains, SYK models, chaotic quantum systems, driven quantum systems, and others [36, 33, 37, 38, 39, 32, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 45, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84]. A detailed explanation on the methods is given in [85]. However, a major challenge in calculating spread and operator complexity is the necessity of classical simulation for constructing the Krylov spaces. In [35], we address this by introducing quantum-mechanically measurable Krylov spaces for the computation of spread complexity and propose Krylov expressivity, a measurable expressivity measure of the phase space dimension. In [34], we extend this research by proposing Krylov observability, a measure to quantify the phase space dimension of various observables measured multiple times. Our results show that Krylov observability effectively captures how well a quantum system can retain and map macroscopic data non-linearly, thus highlighting the importance of Krylov spaces in quantum reservoir computing.
The authors in Ref. [86] first explored the insights that spread complexity can provide in the field of quantum reservoir computing. The authors found that quantum reservoirs with larger mean spread complexity tended to perform better.
In [87] the authors discuss how state estimation in a quantum extreme learning machine can occur beyond the scrambling time, which challenges the common belief that information cannot be retrieved after this point. A similar pattern is observed in quantum reservoir computing, where task performance initially increases and then saturates [88]. We show that fidelity and spread complexity can only explain the initial change in task performance, but fails to explain the saturation in task performance. Our work explains this saturation using the concepts of Krylov expressivity and Krylov observability. The initial state or data is mapped onto the Krylov space, where expressivity and observability increase before reaching a saturation point. Beyond this point, further increases in time do not enhance the system’s expressivity, resulting in stable task performance.
Among all measures, Krylov observability and information processing capacity explain the general trend of task performance best and show almost identical behavior for larger readout dimensions. For small readout dimensions, we show that Krylov observability performs better than information processing capacity in capturing the behavior of the Lorenz task. This is due to the fact that information processing capacity is upper-bounded by the readout dimension, while Krylov observability adapts better to a smaller number of measurements. Our results suggest that Krylov expressivity is not as insightful as Krylov observability in quantum reservoir computing. This is because Krylov expressivity only considers how input states are mapped onto the Krylov space. However, if the initial state is sampled from an encoding unitary, Krylov expressivity can further the understanding of different encoding strategies. Two sub-fields of quantum machine learning with various encoding strategies are variational quantum machine learning and quantum extreme learning [89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113]. Krylov observability, on the other hand, can provide insights into how much of the Krylov space of a variational quantum algorithm can be sampled. Current research on the explainability of quantum machine learning is sparse and primarily employs techniques from classical machine learning in QML [114, 115, 116, 117], while [118] explores the possibilities of designing explanation techniques for parameterized quantum circuits.
This work is organized as follows. Section II introduces quantum reservoir computing, information processing capacity, and the Lorenz task. Section III discusses Krylov spaces in quantum mechanics. Here, we discuss that any time-evolved state lies within a Krylov space and explain measurable Krylov spaces, leading to the introduction of Krylov expressivity. This concept is extended to operator complexity, along with a brief discussion of Krylov observability as proposed in [34]. In Section IV, we demonstrate the limitations of fidelity and spread complexity in explaining task performance due to their oscillatory behavior, even though Lorenz task performance remains mostly constant. We show that Hamiltonians with higher expressivity can exhibit worse performance when Krylov observability is lower. We conclude our work with a discussion of the measures presented.
II Quantum Reservoir Computing
Figure 1: A time series $u_{n}$ is encoded into a qubit of a quantum system, described by a Hamiltonian $H$. The system state is described by the density matrix $\rho_{enc}$ evolving over time $t$. Measurements of some observables $Z_{i}$ at times $m\tau$ are taken to construct the state matrix $\mathbf{S}$. The total number of measurements per observable is $V$ and the clock cycle is defined as $V\tau=T$. In the output layer, the state matrix $\mathbf{S}$ is multiplied by the readout weights $\mathbf{W}^{out}$ to construct the output vector $\mathbf{Y}=\mathbf{S}\mathbf{W}^{out}$.
Quantum reservoir computing makes use of a quantum system as a reservoir for classical or quantum tasks. Fig. 1 shows a sketch of a quantum reservoir scheme and is explained in the following.
  1. 1.
    Initialization: The reservoir is first initialized with an initial time series, in our case of length $N_{\mathrm{In}}=10,000$
  2. 2.
    Input layer (green) and encoding (red): The encoding of the $n$-th input $u_{n}\in\{u_{1},..,u_{N_{u}}\}$ into the first qubit leads to the encoded state
    |$\ket{\Psi_{n}}=\sqrt{\frac{1-u_{n}}{2}}\ket{0}+\sqrt{\frac{1+u_{n}}{2}}\ket{1}.$ (1)
    After encoding the state of the system is described by
[…]
    The overwriting of the first qubit at each time-step results in loss of information about previous time steps. This is called fading memory property and is a crucial requirement for reservoir computing, as reservoirs exhibit a memory-nonlinearity trade-off [1, 2, 119, 120, 121]. In [25] this memory-nonlinear trade-off was shown in quantum reservoir computing.
  3. 3.
    Hidden Layer (blue): The reservoir evolves for a clock cycle $T$ by the unitary evolution
    $U_{R}=\exp(-iHT).$
    A discrete set of observables $\{Z_{i}\}_{i=1..,K}$ is measured to construct the expectation values. The expectation values at time $t$ after the $n$-th input are given by
    $\expectationvalue{Z_{i}(t)}=Tr(Z_{i}e^{-iHt}\rho_{enc}e^{iHt}).$ (3)
    A sketch of the evolution of an observable in regards to different inputs is given above. To increase the readout dimension, we sample each observable $V$ times in each clock cycle $T$, i.e. at times $t_{j}=nT+(j+1)\tau$ with $\tau=T/V$. The sketch Fig. 1 shows a sketch of a quantum reservoir with four virtual nodes $V=4$. This results in $N_{R}=VN_{S}$ readout nodes for each input $u_{n}$, where $N_{S}$ represents the number of observables measured. Writing the readouts for each input into a row, results in the state matrix $\mathbf{S}\in\mathbb{R}^{N_{u},N_{R}}$
  4. 4.
    Output layer (yellow): The state matrix is multiplied with readout weights $\mathbf{W}^{out}$ resulting in the output
    $\mathbf{Y}=\mathbf{S}\mathbf{W}^{out}.$ (4)
  5. 5.
    Training: For training, the data is separated into a initialization set $u_{\mathrm{In}}$, a training set $u_{\mathrm{Tr}}$ and a testing set $u_{\mathrm{Te}}$ of lengths $N_{\mathrm{In}}$, $N_{\mathrm{Tr}}$ and $N_{\mathrm{Te}}$ respectively. A buffer of length $N_{b}$ is added between the sets. We first initialize the system with $N_{\mathrm{In}}$ steps, afterwards the weights are optimized in regards to the training set $u_{\mathrm{Tr}}$, resulting in the trained weights $\mathbf{W}^{out}$. The testing set is used to check if the reservoir can successfully generalize to new data. The weights are optimized to minimize the loss
    $L$ $:=(\mathbf{Y}-\mathbf{Y}^{targ})^{2},$ (5)
    where $\mathbf{Y}^{targ}$ is the target vector. To consider shot noise of the measurement, Gaussian distributed noise $\mathcal{N}\in\mathbb{R}^{N_{u},N_{R}}$ will be added on the state matrix
    $\mathbf{S}$ $\leftarrow\mathbf{S}+\eta\mathcal{N}.$ (6)
    The noise term $\eta\mathcal{N}$ can be used as a regularization parameter when computing the state matrix, this is typically addressed as regularization by noise. The readout weights $\mathbf{W}^{out}$ are then computed by
    $\mathbf{W}^{out}$ $=(\mathbf{S}_{Tr}^{T}\mathbf{S}_{Tr})^{-1}\mathbf{S}_{Tr}^{T}\mathbf{Y}^{targ}.$ (7)
Transverse field Ising Hamiltonian
Four transverse field Ising Hamiltonian with different inter-spin couplings will be used as quantum reservoirs and analyzed in regards to spread complexity, fidelity, Krylov observability, Krylov expressivity and task performance.
${H}_{I\alpha}=\sum_{i=1,j>i}^{N_{S}}J_{ij}X_{i}X_{j}+\sum_{i=1}^{N_{S}}hZ_{i},$ (8)
where $\alpha\in\{1,2,3,4\}$ is the index referencing each of the four Hamiltonian, $N_{S}$ is the number of qubits, $X_{i},Y_{i}$ and $Z_{i}$ are the $\sigma_{x}$, $\sigma_{y}$ and $\sigma_{z}$ Pauli matrices of the $i$th qubit given by
$X_{i},Y_{i},Z_{i}=\Big{(}\bigotimes_{k=1}^{i-1}I_{2}\Big{)}\otimes\sigma_{x,y,z}\otimes\Big{(}\bigotimes_{k=i+1}^{N_{S}}I_{2}\Big{)},$ (9)
and $h=0.5$ is set for all Hamiltonians. The number of pairwise distinct eigenvalues $d$ and the inter-spin couplings $J_{i,j}$ are shown in Table 1.
$H_{I1}$ $H_{I2}$ $H_{I3}$ $H_{I4}$
# of eigenvalues $d$ 9 16 15 16
$J_{1,2}$ 0.5 0.4 0.35 0.35
$J_{1,3}$ 0.5 0.5 0.4 0.4
$J_{1,4}$ 0.5 0.5 0.45 0.45
$J_{2,3}$ 0.5 0.5 0.5 0.5
$J_{2,4}$ 0.5 0.5 0.55 0.55
$J_{3,4}$ 0.5 0.5 0.6 0.65
Table 1: The number of pairwise distinct eigenvalues $d$ and the inter-spin couplings $J_{i,j}$ for the different Ising reservoirs are shown.
Information processing capacity
The information processing capacity (IPC) is a measure of how well a reservoir can generalize prior data onto a set of orthogonal functions [31]. In this study the Legendre polynomials are used. The first order $IPC_{1}$, also called linear IPC, memory capacity or linear short term memory, is a measure of how well the system can remember prior data [122]. Here, the first order Legendre polynomial is used as the target function. The second order $IPC_{2}$ include all target functions consisting of second order Legendre polynomials and products of two first order IPCs with different delay times. Higher order $IPC_{i}$ are defined in a similar fashion, as discussed thoroughly in [1, 25]. The total IPC, or just IPC, is given by the sum of all orders of the IPC:
$\mathrm{IPC}=\sum_{i}IPC_{i}.$ (10)
$\mathrm{IPC}$ is often referred to as memory capacity instead of information processing capacity [123, 124, 125]. In [126, 127], analytical evidence of the influence of eigenvalues on the $\mathrm{IPC}$ in classical reservoir computing is provided. Furthermore it was shown that the $\mathrm{IPC}$ can help in predicting task performance in reservoir computing [128, 129]. The $\mathrm{IPC}$ has been analyzed in regards to quantum reservoir computing in [88, 25].
Lorenz Task
The Lorenz63 system is a commonly used to construct time series prediction benchmark tasks [30]. The dynamics of the system are governed by Eq. 11
$\dot{x}$ $=a(y-x)$
$\dot{y}$ $=x(b-z)-y$
$\dot{z}$ $=xy-cz.$ (11)
With the parameters set to $a=10$, $b=28$, $c=8/3$, the Lorenz63 system exhibits chaotic dynamics. We simulate Eqs. (11) with an integration time step $dt=0.001$. To construct the input and target data for the task, we down sample to a discretized step of $\Delta t=0.02$, which results in the time series $x_{n}=x(n\Delta t)$, $y_{n}=y(n\Delta t)$, and $z_{n}=z(n\Delta t)$. We consider two time series prediction tasks. The first considers the time series $x_{n}$ as input and tries to predict the $p$th future step $x_{n+p}$, referred to as the LXX task. In our study we choose to predict $\Delta t=0.1$ into the future, which results in a $p=5$ step ahead prediction. The second task takes $x_{n}$ as input but tries to predict $z_{n}$. This task is referred to as the LXZ task or cross-prediction task.
Error Measure
The normalized root mean squared error (NRMSE) between output $Y$ of the reservoir and target $Y^{targ}$ is defined as
$\mathrm{NRMSE}=\sqrt{\frac{(\mathbf{Y}-\mathbf{Y}^{\mathrm{targ}})^{2}}{N\mathrm{var}(\mathbf{Y}^{\mathrm{targ}})}}=\sqrt{1-C(\mathbf{Y},\mathbf{Y}^{\mathrm{targ}})}$ (12)
where the capacity or Pearson correlation is defined as
$C(\mathbf{Y},\mathbf{Y}^{\mathrm{targ}})=\frac{\mathrm{cov}(\mathbf{Y},\mathbf{Y}^{\mathrm{targ}})}{\sigma^{2}(\mathbf{Y})\sigma^{2}(\mathbf{Y}^{\mathrm{targ}})}.$ (13)
$\mathrm{cov}(\cdot)$ and $\sigma(\cdot)$ are the covariance and standard deviation, respectively.
Simulation parameters
In Table 2 the simulation parameters for the computation of the results are listed. The information processing capacity is computed up to the fourth order.
value
$N_{\mathrm{In}}$ 10.000
$N_{\mathrm{Tr}}$ 25.000
$N_{\mathrm{Te}}$ 5.000
number of spins $N_{S}$ 4
multiplexing $V$ 15
regularization $\eta$ $10^{-5}$
buffer $N_{b}$ 100
Table 2: QRC and QELM simulation parameters.
III Krylov Spaces in Quantum Mechanics
This chapter introduces Krylov spaces in relation to quantum evolution. We first discuss spread complexity [32], a measure which quantifies the spread over the Krylov basis. Krylov expressivity is defined as a quantum-mechanically measurable quantity that quantifies the effective phase space dimension of evolved quantum states [35]. Next, operator complexity is explored, defining a measure of how operators spread over a Krylov operator basis [33]. Building on this, Krylov observability is introduced as a measure that effectively captures the phase-space dimension of a set of operators ${O_{1},..,O_{K}}$, considering the number of measurements involved in constructing the phase space. In [34], we demonstrate that Krylov observability effectively captures the behavior of the $\mathrm{IPC}$ and the generalization properties of the quantum reservoir.
III.1 Krylov Spaces for State Evolution
The observation of the time-evolution operator being a map onto a Krylov space was first discussed in [33, 32]. There, the proof for Eq. 17 was given, which we will be going over briefly.
Spread Complexity
The Schrödinger equation for the time-independent Hamiltonian $H$ with initial condition |$\ket{\Psi(0)}$ is given by:
$\partial_{t}\ket{\Psi(t)}=-iH\ket{\Psi(t)}$
|$\ket{\Psi(0)}:=\ket{\Psi_{0}}.$ (14)
The solution to this equation is
|$\ket{\Psi(t)}=e^{-iHt}\ket{\Psi_{0}}=\sum_{k=0}^{\infty}\frac{(-iHt)^{k}}{k!}\ket{\Psi_{0}},$ (15)
where the series representation of the matrix exponent was used. Defining the linear function $f(\ket{\Psi_{0})}:=-iH$ leads to
|$\ket{\Psi(t)}$ $=e^{-iHt}\ket{\Psi_{0}}$
$=\sum_{k=0}^{\infty}(-iH)^{k}\frac{t^{k}}{k!}\ket{\Psi_{0}}=\sum_{k=0}^{\infty}f^{k}(\ket{\Psi_{0}})\frac{t^{k}}{k!}.$ (16)
For any time $t\in\mathbb{R}$ it holds that
|$\ket{\Psi(t)}\in\mathrm{Span}\{f^{0}(\ket{\Psi_{0}}),f^{1}(\ket{\Psi_{0}})t,f^{2}(\ket{\Psi_{0}})\frac{t^{2}}{2!},\ldots\}$
Since $t^{k}/k!>0$ for $t>0$ it follows that
|$\ket{\Psi(t)}\in\mathrm{Span}\{f^{0}(\ket{\Psi_{0}}),f^{1}(\ket{\Psi_{0}}),\ldots\}:=K_{\infty}.$
$f$ being a linear function and considering the Krylov space property implies that there exists a $m\leq N$, such that
$K_{\infty}$ $=\mathrm{Span}\{f^{0}(\ket{\Psi_{0}}),f^{1}(\ket{\Psi_{0}}),f^{2}(\ket{\Psi_{0}}),\ldots\}$
$=\mathrm{Span}\{f^{0}(\ket{\Psi_{0}}),f^{1}(\ket{\Psi_{0}}),\ldots\,f^{m-1}(\ket{\Psi_{0}})$
$=\mathrm{K}_{m}$ (17)
$m$ is called the grade of |$\ket{\Psi_{0}}$ in regards to $f:=-iH$ [33, 32]. Building upon this [32] proposed the spread complexity, which is a measure of how the state spreads over the Krylov basis of $\mathrm{K}_{m}$. This is achieved by orthonormalization through the Lanczos algorithm of the vectors $f^{j}(\ket{\Psi_{0}})$, which results in
$\mathrm{K}_{m}=\mathrm{Span}\{\ket{k_{0}},..,\ket{k_{m-1}}\}.$ (18)
The time evolved state is reconstructed through the basis representation |$\ket{k_{i}}$ as
|$\ket{\Psi(t)}=\sum_{n=0}^{m-1}\bra{k_{n}}\ket{\Psi(t)}\ket{k_{n}}.$ (19)
With $\alpha_{n}(t)=\bra{k_{n}}\ket{\Psi(t)}$, the spread complexity is defined as
$\mathcal{K}_{S}(t)=\sum_{n=0}^{m-1}(n+1)\absolutevalue{\alpha_{n}(t)}^{2}.$ (20)
Spread complexity characterizes how an initial state evolves over a basis constructed from different powers of the Hamiltonian. This metric provides valuable insights and helps in the understanding of the time evolution operator, but faces some challenges as an expressivity measure for quantum machine learning and quantum reservoir computing. First, the requirement for knowledge of the system’s Hamiltonian poses a challenge because the quantum machine learning network is typically not described by a Hamiltonian but by a series of quantum circuits. Additionally, the different powers of the Hamiltonian must be computed classically, which becomes increasingly difficult as the number of qubits grows. To address these challenges, we demonstrated that time-evolved states can be used to construct a basis from which the complexity measure can be defined, exhibiting the same behavior in the researched systems [35].
Krylov Expressivity
In [35] we demonstrated that instead of using the Krylov space
$\mathrm{K}_{m}=\mathrm{Span}\Big{(}\ket{\Psi_{0}},H\ket{\Psi_{0}},\ldots,H^{m-1}\ket{\Psi_{0}}\Big{)},$
a set of time-evolved states can be employed to construct the space
$\mathrm{G}_{m}:=\mathrm{Span}\Big{(}e^{-iHt_{0}}\ket{\Psi_{0}},\ldots,e^{-iHt_{m-1}}\ket{\Psi_{0}}\Big{)}.$
The basis of $\mathrm{G}_{m}$ are evolved states and therefore quantum-mechanically measurable. The global phase does not change the space dimension and detailed proofs are given in [35]. In [35] we showed that the grade $m$ of the Krylov space $\mathrm{K}_{m}$ is equal to the number of pairwise distinct eigenvalues of the Hamiltonian $d$, i.e., $m=d$. Building on this, Krylov expressivity is defined, which is upper bounded by the grade $m$.
The computation of Krylov expressivity requires the knowledge of the grade $m$. For this, time evolved states |$\ket{g_{i}}=\exp(-iHt_{i})\ket{\Psi_{0}}$ are computed, where some $T_{K}$ is picked and $t_{i}=(i+1)T_{K}/(N+1)$. For exactly the grade $m$, it will hold that |$\ket{g_{m}}\in\mathrm{Span}(\ket{g_{1}},\ket{g_{2}},..,\ket{g_{m-1}})=\mathrm{G}_{m}$. We then simulate the time evolved state until some time $T$ and calculate the sampling times $\tau_{i}=(i+1)T/m$ with $i=0,..,m-1$. For small $T$ it is expected that the vectors |$\ket{g_{i}}$ are close to each other, i.e. for all $i$, |$\ket{g_{i}}\approx\ket{g_{i+1}}$ would hold. Analytically speaking |$\ket{g_{i}}$ are independent, but numerically almost equal. For the computation of the Krylov expressivity, a measure of similarity between two time-evolved states |$\ket{g_{i}}$ and |$\ket{g_{i+1}}$ is required. Here, we proposed the computation of the fidelity for pure states $\mathrm{F}(\ket{g_{i}},\ket{g_{i+1}})$ but other measures can also be utilized.
[…]
The Krylov expressivity $\mathcal{E}_{K}$ for $\lambda$, where $\lambda\in[0,1]$ holds, is given by
$m_{{eff}_{i}}$ =$=\begin{cases}1&\mathrm{if~{}}\lambda_{i}<\lambda\\ 1-\frac{1}{1-\lambda}\cdot(\lambda_{i}-\lambda)&\mathrm{if~{}}\lambda_{i}\geq\lambda\end{cases}$
$\mathcal{E}_{K}$ $=1+\sum_{i=1}^{m-1}m_{{eff}_{i}}.$ (22)
The first vector |$\ket{g_{1}}$ adds a dimension of $1$. For $\lambda_{i}<\lambda$ we say that the two vectors are linearly independent and the Krylov expressivity is increased by one. In the region $\lambda_{i}\geq\lambda$, the Krylov expressivity $m_{{eff}_{i}}$ is interpolated to consider the difference between analytical independence and the independence that is needed for computation. We will explore the Krylov expressivity $\mathcal{E}_{K}$ in dependence of the time scale $T$ and which influence this has on the performance in quantum reservoir computing.
III.2 Krylov Spaces for Operator Evolution
It is known, that where spread complexity tries to quantify the spread of states, Operator complexity quantifies the spread of time-evolved operators [33] and is defined in the following paragraph.
Operator complexity
The evolution of operators is governed by
$\partial_{t}O(t)=i[H,O(t)].$ (23)
The solution to this equation is
$O(t)$ $=e^{iHt}Oe^{-iHt}=\sum_{k=0}^{\infty}\frac{(it)^{k}}{k!}\mathcal{L}^{k}(O),$ (24)
where $\mathcal{L}(O)=HO-OH$ is the Liouvillian superoperator and linear in $O$, i.e.
$\mathcal{L}(\lambda O_{1}+O_{2})$ $=H(\lambda O_{1}+O_{2})-(\lambda O_{1}+O_{2})H$
$=\lambda HO_{1}-\lambda HO_{1}+HO_{2}-HO_{2}$
$=\lambda\mathcal{L}(O_{1})+\mathcal{L}(O_{2})$ (25)
Any time evolved operator $O(t)$ has therefore to be in the span of the powers of the Liouvillian $\mathcal{L}^{n}$
$O(t)\in\mathrm{Span}\{\mathcal{L}^{0}(O),\mathcal{L}^{1}(O),\mathcal{L}^{2}(O),..\}.$ (26)
By using the linear property of $\mathcal{L}$ and the Krylov space property, authors of [33] showed that there exists a $M\in\mathbb{N}$ such that
$O(t)\in\mathrm{Span}\{\mathcal{L}^{0}(O),\mathcal{L}^{1}(O),..,\mathcal{L}^{M-1}(O),..\}=\mathcal{L}_{M}$ (27)
holds true. Similar to the Krylov state complexity the authors perform an orthonormalization procedure to construct the space
$\mathcal{L}_{M}$ $:=\mathrm{Span}\{\mathcal{L}^{0}(O),\mathcal{L}^{1}(O),..,\mathcal{L}^{M-1}(O)\}$
$=\mathrm{Span}\{\mathcal{W}_{0},\mathcal{W}_{1},..,\mathcal{W}_{M-1}\},$ (28)
where $\{\mathcal{W}_{i}\}_{i=..,M-1}$ is the orthonormal basis constructed through the Lanczos algorithm. In the next step the time evolved operator is expressed in the Krylov basis as
$O(t)=\sum_{n=0}^{M-1}i^{n}\beta_{n}(t)\mathcal{W}_{n},$ (29)
where $\beta_{n}(t)=(O(t),\mathcal{W}_{n})$ is a scalar product defined on the operator space. Operator complexity $\mathcal{K}_{O}$ is then defined as
$\mathcal{K}_{O}(t)=\sum_{n=0}^{M-1}(n+1)\absolutevalue{\beta_{n}(t)}^{2}.$ (30)
Krylov Observability
One aspect that operator complexity lacks is a way to define an effective space dimension and a method to consider multiple operators. This is especially crucial for quantum machine learning, where a set of observables is measured, and quantum reservoir computing, where multiplexing is used to increase the readout dimension. In [34] we show that instead of using the Krylov space of the different powers of the Liouvillian $\mathcal{L}^{k}$
$O(t)\in\mathcal{L}_{M}:=\mathrm{Span}\{\mathcal{L}^{0}(O),\mathcal{L}^{1}(O),..,\mathcal{L}^{M-1}(O)\}$
, we can use time-evolved observables to construct an equivalent space.
$O(t)\in\mathcal{F}_{M}=\mathrm{Span}\{\tilde{O}(t_{0}),\tilde{O}(t_{1}),..,\tilde{O}(t_{M-1})\}=\mathcal{L}_{M}$ (31)
Building on this, we present an algorithm (Appendix A) that enables the computation of a minimal space $\mathcal{F}$ such that any time-evolved operator $O_{i}(t)\in\mathcal{F}$ holds for all $O_{i}\in\{O_{1},O_{2},\dots,O_{K}\}$ and all $t$. This algorithm further returns the contribution of each operator $O_{i}$ as a space $\mathcal{F}_{i}$ with the following properties
$\mathcal{F}=\bigcup_{k=1}^{K}\mathcal{F}_{k}=\bigcup_{k=1}^{K}\tilde{\mathcal{F}}_{k}~{},~{}~{}\mathrm{dim}\left(\bigcup_{j=1}^{l}\tilde{\mathcal{F}}_{j}\right)=\sum_{j=1}^{l}\mathrm{dim}(\tilde{\mathcal{F}}_{j}).$
The computation of Krylov observability considers $K$ observables $O_{1},\dots,O_{K}$[34]. For each observable $O_{k}$, the spaces ${\mathcal{F}}_{k}$ are computed with $\mathrm{dim}({\mathcal{F}}_{k})=M_{k}$. With $R_{k}=\mathrm{min}(V,M_{k})$ and $\tau_{k}=T/M_{k}$, we can then define
$\mathcal{G}_{k}=\{O_{k}(\tau_{1}),O_{k}(\tau_{2}),\dots,O_{k}(\tau_{M_{k}})\},$
where $V$ is the number of measurements. The observability of the $k$-th observable $O_{k}$ is defined as
$\kappa_{k}(T)=1+\sum_{j=1}^{R_{k}-1}(1-\mathrm{F}(O_{k}(\tau_{j}),O_{k}(\tau_{j+1})))$ (32)
, where $\mathrm{F}$ is the normalized fidelity between the two time-evolved operators given by
$\mathrm{F}(A,B)=\absolutevalue{\mathrm{Tr}\Big{(}\frac{A^{\dagger}B}{\norm{A}\norm{B}}\Big{)}}.$ | . (33)
The Krylov Observability $\mathcal{O}_{K}(T)$ of $V$ multiplexed observables $O_{1},\dots,O_{K}$ is defined as
$\mathcal{O}_{K}(T)=\sum_{k=1}^{K}\kappa_{k}(T).$ (34)
IV Results
Figure 2: Fidelity $\mathrm{F}$ (black) and spread complexity $\mathcal{K}_{S}$ (red) in (a), Krylov observability $\mathcal{O}_{K}$ (red) and Krylov expressivity $\mathcal{E}_{K}$ in (b) and the Lorenz task performance for the five-step ahead prediction ($\Delta t=0.1$) of the $x$-variable (green) and the cross prediction of the $z$- variable (blue) in (c) in dependence of the clock cycle $T$, when measuring one site of the $H_{I4}$ Hamiltonian. The bold line in (c) is the fit to the $\mathrm{NRMSE}$ to better visualize the saturation in task performance.
The aim of this work is to utilize and compare expressivity measures to deepen the understanding of quantum reservoir computing. To begin, we simulate spread complexity and fidelity to investigate whether task performance can be explained by these measures. Fig. 2.a illustrates fidelity and spread complexity averaged over twenty random starting states. We see that spread complexity $\mathcal{K}_{S}$ increases, while fidelity $F$ decreases with larger clock cycles. Additionally, we observe an anti-proportional relationship between the two measures. When fidelity reaches a minimum at $T\approx 36$, spread complexity exhibits a maximum, and when fidelity increases, spread complexity decreases, and vice versa.
In Fig. 2.c, the $\mathrm{NRMSE}$ values of the Lorenz tasks for the five-step prediction of the $x$ variable (green) and the cross-prediction of the $z$ variable (blue) are presented. We observe a reduction in error, followed by a saturation in task performance. The initial decrease in $\mathrm{NRMSE}$ can be explained using fidelity, which captures how far the state evolves from the initial state. This however only tells us that the system has to evolve, to perform any operations as at $T=0$ the reservoir evolution is simply the identity. Meanwhile, spread complexity quantifies how the time-evolved state spreads over time in the Krylov basis. Higher spread complexity therefore implies a broader spread within the Krylov basis at the corresponding clock cycle. However, neither fidelity nor spread complexity alone can explain the saturation of $\mathrm{NRMSE}$. The oscillatory behavior observed in fidelity and spread complexity (Fig. 2.a) is not reflected in task performance (Fig. 2.c). Fig. 2.b shows Krylov expressivity $\mathcal{E}_{K}$ and Krylov observability $\mathcal{O}_{K}$. Both measures show an increase followed by saturation, around which they oscillate, similar to the task performance in Fig. 2.c. This shows that Krylov expressivity and Krylov observability capture the saturation in task performance, while only limited insights can be gained from fidelity and spread complexity.
Figure 3: The first row shows the information processing capacity $\mathrm{IPC}$ (black), the Krylov expressivity $\mathcal{E}_{K}$ (red) and the Krylov observability $\mathcal{O}_{K}$ (purple) and the second row shows the Lorenz task performance for the five-step ahead prediction ($\Delta t=0.1$) of the $x$-variable (green) and the cross prediction of the $z$- variable (blue) in dependence of the clock cycle $T$ for the different quantum reservoirs a) $H_{I1}$, b) $H_{I2}$,c) $H_{I3}$ and d) $H_{I4}$. The state matrix is constructed by measuring the $Z_{1}$ observable $V=30$ times.
Fig. 3 shows the results when only the first observable $Z_{1}$ is measured, while in Fig. 4, measurements of the observables $Z_{1}$, $Z_{2}$, $Z_{3}$, and $Z_{4}$ are considered. In all cases, the observables are multiplexed $V=30$ times. A summary of the results is provided in Appendix B. The Krylov expressivity $\mathcal{E}_{K}$ of each Hamiltonian $H_{I1}$, $H_{I2}$, $H_{I3}$, and $H_{I4}$ is upper-bounded by $9$, $16$, $15$, and $16$, respectively. This upper bound corresponds to the grade of the Krylov space or the number of pairwise distinct eigenvalues. Fig. 3.a shows the result for $H_{I1}$. We observe that $\mathrm{NRMSE}$ of the Lorenz task, $\mathrm{IPC}$, and $\mathcal{O}_{K}$ all saturate around $T=12$, while $\mathcal{E}_{K}$ reaches saturation at $T=6$. Increasing the complexity of the Hamiltonian slightly by modifying one of the inter-spin couplings $J_{i.j}$ results in $H_{I2}$. This Hamiltonian reaches a higher $\mathrm{IPC}$, Krylov expressivity $\mathcal{E}_{K}$ and Krylov observability $\mathcal{O}_{K}$. Furthermore the Lorenz task errors are slightly decreased. When using pairwise distinct inter-spin couplings $J_{i,j}$ with equal spacing, the resulting Hamiltonian is $H_{I3}$ (Fig. 3.c. Here, $\mathcal{E}_{K}$, $\mathrm{IPC}$, and $\mathcal{O}_{K}$ each saturate at $T_{sat}\approx 8$, $13$, and $20$, respectively. The Lorenz task performance, however, reaches saturation at $T_{sat}\approx 18$, which best aligns with the behavior of Krylov observability $\mathcal{O}_{K}$.
For $H_{I1}$ and $H_{I2}$, $\mathrm{IPC}$ and $\mathcal{O}_{K}$ show similar behavior, but a discrepancy appears in $H_{I3}$. This occurs because $\mathrm{IPC}$ is upper-bounded by the readout dimension of $N_{R}=30$, to which the $\mathrm{IPC}$ then saturates $\mathrm{IPC}\leq N_{R}$. In this case, Krylov observability $\mathcal{O}_{K}$ best explains the behavior of task performance. The worse task performance of $H_{I1}$ compared to the other Hamiltonians is explained by the fact that all metrics $\mathrm{IPC}$, $\mathcal{E}_{K}$, and $\mathcal{O}_{K}$ are all smaller.
When comparing $H_{I3}$ and $H_{I2}$ we note that $\mathcal{E}_{K}(H_{I3})<\mathcal{E}_{K}(H_{I2})$. At the same time, we observe that the task performance for the cross-prediction task is better for $H_{I3}$, while it is worse for the prediction of the $x$ variable. To further investigate this discrepancy, we introduce $H_{I4}$. This Hamiltonian is nearly identical to $H_{I3}$, differing only in one inter-spin coupling, which results in a Krylov expressivity equal to that of $H_{I2}$. Despite this adjustment, the $\mathrm{IPC}$, $\mathcal{O}_{K}$, $\mathcal{E}_{K}$, and Lorenz task errors for $H_{I4}$ remain nearly identical to those of $H_{I3}$.
Figure 4: Same as in Fig. 3, except that the state matrix is constructed by measuring the $Z_{1}$, $Z_{2}$, $Z_{3}$, and $Z_{4}$ observables $V=30$ times.
To discuss this further, we consider a typical quantum reservoir where each observable is measured, as shown in Fig. 4. This results in a readout dimension of $N_{R}=120$. In this setup, $H_{I1}$ again exhibits largest errors. While the Lorenz task performance remains unchanged, the information processing capacity increases from $\mathrm{IPC}=20$ (Fig. 3) to $\mathrm{IPC}=40$ in Fig. 4.a. For $H_{I2}$, $H_{I3}$, and $H_{I4}$ (Fig. 4.b-d, we observe that $\mathrm{IPC}$ and $\mathcal{O}_{K}$ both perform similarly well in explaining the saturation of the $\mathrm{NRMSE}$ of the Lorenz task.
The improved performance of $\mathrm{IPC}$ in Fig. 4 compared to Fig. 3 is due to the system being sampled with $N_{R}=120$ readout nodes. In this case the $\mathrm{IPC}$ is upper-bounded by $\mathrm{IPC}\leq N_{R}=120$, which is sufficiently large to not observe the upper bound by the readout dimension.
$Z_{1}$
$x\rightarrow x$,
$\mathrm{NRMSE}_{sat}$
all sites
$x\rightarrow x$
$\mathrm{NRMSE}_{sat}$
$Z_{1}$
$x\rightarrow z$,
$\mathrm{NRMSE}_{sat}$
all sites
$x\rightarrow z$
$\mathrm{NRMSE}_{sat}$
$H_{1}$ 0.08 0.08 0.3 0.3
$H_{2}$ 0.04 0.03 0.2 0.15
$H_{3}$ 0.06 0.03 0.16 0.08
$H_{4}$ 0.06 0.03 0.16 0.08
Table 3: $\mathrm{NRMSE}$ for the Lorenz task of the four Hamiltonians $H_{I1},~{}H_{I2},~{}H_{I3}$ and $H_{I4}$, when measuring one site $Z_{1}$ (row one and three) and when measuring all sites (row two and four).
Table 3 shows the $\mathrm{NRMSE}$ of the Lorenz task at saturation, when only $Z_{1}$ is measured and when all four sites are measured. For $H_{I1}$, we observe that the number of sites measured does not change task behavior. For $H_{I2}$, we observe a slightly smaller $\mathrm{NRMSE}$ when all sites are measured. $H_{I3}$ and $H_{I4}$ exhibit identical task performance, with the error of both tasks halved when more observables are measured.
To discuss the different behavior of expressivity and observability of $H_{I2}$ and $H_{I3}$, we take a closer look at the task performance when all sites are measured (Fig. 4). We observe that the cross-prediction task for $H_{I3}$ has a $\mathrm{NRMSE}=0.08$, while for $H_{I2}$, the error is $\mathrm{NRMSE}=0.15$. The Krylov expressivity for $H_{I2}$, with $\mathcal{E}_{K}(H_{I2})=16$, is larger than for $H_{I3}$, which has $\mathcal{E}_{K}(H_{I3})=15$. This seems counterintuitive, since $H_{I2}$ maps input data onto a larger Krylov space, and the amount of information that can be extracted should also be greater. However, even though the data is mapped onto a larger space, the information is extracted through a set of measured observables. In this specific case, we observe that the number of independent measurements for $H_{I2}$ is smaller than for $H_{I3}$, which is quantified by the smaller Krylov observability $\mathcal{O}_{K}(H_{I2})=60$ compared to $\mathcal{O}_{K}(H_{I3})=85$.
This behavior is likely due to commutation between the observables and a portion of the subspace mapped by the quantum reservoir in $H_{I2}$. By modifying the inter-spin couplings to be pairwise distinct, the observables gain access to more information about the quantum state, leading to increased observability.
To further probe this behavior, the $\mathrm{IPC}$ and $\mathcal{O}_{K}$ are computed for various clock cycles, virtual nodes, and numbers of readout observables for $H_{I2}$ (Fig. 5) and $H_{I3}$ (Fig. 6). In the first and second rows, only the first site and the first two sites are measured for the construction of the state matrix, respectively, while the third row shows the results when all sites are measured.
For $H_{I2}$ (Fig. 5), we observe $\mathrm{IPC}<85$ and Krylov observability $\mathcal{O}_{K}<85$. In contrast, $H_{I3}$ shows a maximum of $105$ (Fig. 6), indicating that the larger $\mathrm{IPC}$ is due to the increased Krylov observability $\mathcal{O}_{K}$.
Furthermore, an increase in the number of observables measured shifts the best performance from larger clock cycles and a larger number of measurements to smaller clock cycles and fewer measurements. Each plot includes the Pearson correlation coefficient $P_{C}$ (Eq. 13) between $\mathrm{IPC}$ and $\mathcal{O}_{K}$. For $H_{I2}$, the correlation coefficients are $P_{C}=0.94$ when one or two observables are measured, and $P_{C}=0.96$ when all observables are considered. This indicates a strong and consistent correlation across all configurations. The corresponding correlation coefficients for $H_{I3}$ are $P_{C}=0.91$ for one observable and $P_{C}=0.95$ for both two and four observables.
Figure 5: Observability $\mathcal{O}_{K}$ (right column) and $\mathrm{IPC}$ (left column) are color-coded from 0 to 85, depending on multiplexing $V$ and clock cycle $T$ for $H_{I2}$. The Pearson correlation factor $P_{C}$ between the two images in each row is calculated, indicating an almost identical behavior between $\mathcal{O}_{K}$ and $\mathrm{IPC}$.
Figure 6: Same as in Fig. 6 for $H_{I3}$.
In a final experiment, we compute the $\mathrm{IPC}$ and $\mathcal{O}_{K}$ for an Ising Hamiltonian with five sites and pairwise distinct couplings. The results are presented in Fig. 7, where one site (first row), three sites (second row), and all five sites (third row) are measured in the $z$-direction. Simulations are performed up to $V=220$ virtual nodes and clock cycles up to $T=200$. Computing the $\mathrm{IPC}$ for each parameter configuration requires hundreds of hours for the five-site system and tens of hours for the four-site system, which limits our ability to simulate even larger systems. Once again, we observe almost identical behavior between $\mathrm{IPC}$ and $\mathcal{O}_{K}$, as indicated by a correlation factor of $P_{C}=0.97$.
Figure 7: Same as in Fig. 6 for for a reservoir with five sites $N_{S}=5$.
V Discussion and Conclusion
This work investigates Krylov-based information measures to explain and predict task performance in quantum reservoir computing, using the Lorenz63 system for chaotic time-series prediction. Initially, fidelity and spread complexity are computed, but they fail to explain task performance saturation Fig. 2.
We then extended our research to Krylov expressivity $\mathcal{E}_{K}$ [35] and Krylov observability $\mathcal{O}_{K}$ [34] and show that both measures show saturation for larger clock cycles. The measures are compared to the information processing capacity $\mathrm{IPC}$, which captures how well the system can retain and map data non-linearly. We test four quantum reservoirs in an undersampled regime, where Krylov observability $\mathcal{O}_{K}$ outperforms $\mathrm{IPC}$ in explaining task performance. In this case, $\mathrm{IPC}$ reaches a maximum, while the error for the Lorenz tasks continues to decrease with larger clock cycles $T$.
The quantum reservoirs are then simulated when all sites are measured, as shown in Fig. 4. The first observation from this experiment is that $\mathrm{IPC}$ and $\mathcal{O}_{K}$ exhibit almost identical behavior. Another significant result is the superior Lorenz task performance and larger $\mathrm{IPC}$ of $H_{I3}$ compared to $H_{I2}$, despite $H_{I3}$ having a smaller Krylov expressivity. While Krylov expressivity provides insights into the space onto which the data is mapped, the amount of information extracted from that space is smaller for $H_{I2}$, as evidenced by the smaller Krylov observability $\mathcal{O}_{K}$ of $H_{I2}$ compared to $H_{I3}$. To confirm that our measures are not dependent on system size, we computed the information processing capacity and Krylov observability for a five-qubit system, as shown in Fig. 7, achieving correlation factors of $P_{C}\geq 0.97$. Since Krylov expressivity considers how input states are mapped onto the Krylov space, information about the input encoding can be gained. In quantum reservoir computing, this might not be of much importance, because the state evolves over time. In quantum machine learning, however, where input encoding is one of the main parts to be optimized, Krylov expressivity can be used to effectively understand and compare various encoding strategies. One approach would be to encode the input data $x_{i}$ through a unitary circuit $U_{E}(x_{i})$ into the quantum reservoir or quantum machine learning network. The initial states in the computation of Krylov expressivity can then be sampled from the set of encoded states, i.e., |$\ket{x}\in\{U_{E}(x_{i})\ket{s}\mid x_{i}\in\mathcal{X}\}$, where $\mathcal{X}$ is the set of inputs. The utility of Krylov observability might be used to further gain understanding in quantum machine learning, such as barren plateaus or as a quantum-mechanical information measure that can be utilized in the understanding of quantum dynamics, further advancing knowledge in chaos, decoherence, or thermalization.
Appendix A Algorithm for the construction of the spaces
Consider the basis $\mathcal{F}^{B}$ and $\mathcal{F}^{B}_{i}$ of the spaces $\mathcal{F}$ and $\mathcal{F}_{i}$, i.e., $\mathcal{F}=\mathrm{Span}(\mathcal{F}^{B})$ and $\mathcal{F}_{i}=\mathrm{Span}(\mathcal{F}^{B}_{i})$. Then Algorithm 1 constructs spaces such that the properties given in Section III.2 are defined.
Algorithm 1 Construction of Observability Spaces
1:$\mathcal{I}_{O}=\{1,2,\dots,K\}$
2:$\mathcal{T}=\{t_{1},t_{2},\dots,t_{R}\}$
3:$\mathcal{F}^{(B)}=\emptyset$
4:~$\tilde{\mathcal{F}}_{1}^{(B)},\tilde{\mathcal{F}}_{2}^{(B)},\dots,\tilde{\mathcal{F}}_{K}^{(B)}=\emptyset,\emptyset,\dots,\emptyset$
5:while $t_{j}\in\mathcal{T}$ do
6:     for $k\in\mathcal{I}_{O}$ do
7:         if $O_{k}(t_{j})\notin\mathcal{F}$ then
8:              $\mathcal{F}^{(B)}=\mathcal{F}^{(B)}\cup O_{k}(t_{j})$
9:              ~$\tilde{\mathcal{F}}_{k}^{(B)}=\tilde{\mathcal{F}}_{k}^{(B)}\cup O_{k}(t_{j})$
10:         end if
11:     end for
12:end while
13:$\mathcal{F}^{(B)}\leftarrow\mathrm{ONB}(\mathcal{F}^{(B)})$
Appendix B Saturation
Below table gives information when each measure exhibits a saturation point.
$\mathcal{E}_{K}$
$T_{sat}$
$\mathcal{O}_{K}$
$T_{sat}$
$\mathrm{IPC}$
$T_{sat}$
$\mathrm{IPC}$
saturation point
$x\rightarrow x$
$T_{sat}$
$x\rightarrow z$
$T_{sat}$
$x\rightarrow x$,
$\mathrm{NRMSE}_{sat}$
$x\rightarrow z$
$\mathrm{NRMSE}_{sat}$
$H_{1}$, ($Z_{1}$) 6 12 12 20 12 12 0.08 0.3
$H_{1}$, ($Z_{1},Z_{2},Z_{3},Z_{4}$) 6 10 34 40 12 12 0.08 0.3
$H_{2}$, ($Z_{1}$) 8 17 15 30 18 18 0.04 0.2
$H_{2}$, ($Z_{1},Z_{2},Z_{3},Z_{4}$) 8 20 22 60 18 18 0.03 0.15
$H_{3}$, ($Z_{1}$) 8 20 13 30 18 18 0.06 0.16
$H_{3}$, ($Z_{1},Z_{2},Z_{3},Z_{4}$) 8 20 20 87 20 20 0.03 0.08
$H_{4}$, ($Z_{1}$) 8 20 13 30 18 18 0.06 0.16
$H_{4}$, ($Z_{1},Z_{2},Z_{3},Z_{4}$) 8 20 20 87 20 20 0.03 0.08
Table 4: Information about the saturation behavior, saturation time and saturation points of the different Hamiltonians.